OpenStudy (anonymous):

A truck traveling at 25m/s puts its brakes on for 4s.This produces a retardation of 2m/s^2.What does the truck's velocity drop to? Please show math(if you can)? Answer: 17m/s

OpenStudy (anonymous):

v = u + at v = 25 - 2*4 v = 17 m/s